Practice Question and Answer

Q:

Complete the series choosing the missing number: 

1, 64, 343, 1000, _____

2027 0

  • 1
    4739
    Correct
    Wrong
  • 2
    6391
    Correct
    Wrong
  • 3
    2197
    Correct
    Wrong
  • 4
    None of the above
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "2197"

Q: Select the most appropriate option to substitute the underlined segment in the given sentence. If there is no need to substitute it, select 'No substitution'.

He waits to you for about two hours.

2027 0

  • 1
    has been waiting to you
    Correct
    Wrong
  • 2
    has waiting for you
    Correct
    Wrong
  • 3
    has been waiting for you
    Correct
    Wrong
  • 4
    No substitution
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 4. "No substitution"

Q:

The shadow of a tower, when the angle of elevation of the sun is 60° is found to be 15 m shorter than when it is 45°. Then height of the tower is 

2027 0

  • 1
    41.5 m
    Correct
    Wrong
  • 2
    35.5 m
    Correct
    Wrong
  • 3
    26.5m
    Correct
    Wrong
  • 4
    20.5m
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "35.5 m "

Q:

मुहावरे और अर्थ का कौनसा विकल्प सही नहीं हैं?

2027 0

  • 1
    लोहा मानना-श्रेष्ठता स्वीकार करना
    Correct
    Wrong
  • 2
    लोहा लेना-मुकाबला करना
    Correct
    Wrong
  • 3
    लल्लो-चप्पो करना-खुशामद करना
    Correct
    Wrong
  • 4
    लकीर का फकीर होना-सही रास्ते पर चलना
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 4. "लकीर का फकीर होना-सही रास्ते पर चलना"

Q: Can you solve the Math Puzzle ? 2027 0

  • 1
    12
    Correct
    Wrong
  • 2
    15
    Correct
    Wrong
  • 3
    17
    Correct
    Wrong
  • 4
    19
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "15"
Explanation :

Answer: B) 15 Explanation: The puzzle's logic is 1st figure : (9-7) + (8-5) = 2 + 3 = 5 2nd figure : (15-6) + (11-2) = 9 + 9 = 18 similarly, 3rd figure : (9-6) + (17-5) = 3 + 12 = 15.

Q: In a box, there are four marbles of white color and five marbles of black color. Two marbles are chosen randomly. What is the probability that both are of the same color? 2027 0

  • 1
    2/9
    Correct
    Wrong
  • 2
    5/9
    Correct
    Wrong
  • 3
    4/9
    Correct
    Wrong
  • 4
    0
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "4/9"
Explanation :

Answer: C) 4/9 Explanation: Number of white marbles = 4 Number of Black marbles = 5 Total number of marbles = 9 Number of ways, two marbles picked randomly = 9C2 Now, the required probability of picked marbles are to be of same color = 4C2/9C2  +  5C2/9C2 = 1/6 + 5/18 = 4/9.

Q:

Complete the series.
TEA, UFB, ?, WHD

2027 0

  • 1
    VGC
    Correct
    Wrong
  • 2
    XIC
    Correct
    Wrong
  • 3
    VJD
    Correct
    Wrong
  • 4
    XIZ
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "VGC"

Q:

Hubli railway station in Karnataka has been renamed to?

2027 0

  • 1
    Nagpur Railway Station
    Correct
    Wrong
  • 2
    Jaipur Railway Station
    Correct
    Wrong
  • 3
    Siddharudha Railway Station
    Correct
    Wrong
  • 4
    Siddharudha Swamiji Railway Station
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 4. "Siddharudha Swamiji Railway Station"

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully